OpenStudy (anonymous):

Every student in the 6th grade at PS 500 must read exactly three of the principal's four favorite books. If 34 read Math for Fun, 36 read Everyone Loves Math, 38 read Cool Math, and 12 read Math is for Me, how many students are in the 6th grade at PS 500? (A) 30 (B) 40 (C) 60 (D) 90 (E) 120

OpenStudy (decarr432):

Welcome to open study is this all thats given?

OpenStudy (decarr432):

I believe I know the answer

OpenStudy (anonymous):

Yes, this is all that is given

OpenStudy (decarr432):

I do believe you just have to add all numbers and there is your answer or You might have to find how many students there are based on each one reading three books

OpenStudy (decarr432):

I'm not sure tho

OpenStudy (anonymous):

Adding everything will be really easy. There is more steps involved.

OpenStudy (decarr432):

I know if you add them all together you get (E)

OpenStudy (anonymous):

FYI the answer is B but I do not know how to solve this problem

OpenStudy (decarr432):

Okay so take all the students \[34+36+38+12= 120 \] \[120\div3=40\]

OpenStudy (decarr432):

Then just divide by three because it says each student has to read three of the four

OpenStudy (anonymous):

I did not think of that. It so simple

OpenStudy (decarr432):

Math can be simple some times

OpenStudy (anonymous):

Thank you for your help

OpenStudy (decarr432):

No problem if you want to you can fan+medal it helps me and it lets me know when you ask a question

OpenStudy (xapproachesinfinity):

the answer is incorrect

OpenStudy (anonymous):

xapproachesinfinity How

OpenStudy (xapproachesinfinity):

why did you divide by 3 can you explain why?

OpenStudy (retireed):

I wrote four equation for the four unknowns, but I have not been able to solve them, while on laptop. Might need to get some paper or you can try them if you want to.

OpenStudy (anonymous):

xapproachesinfinity How do I know? Ask Decarr432

OpenStudy (xapproachesinfinity):

I'm asking him not you? unless he explains why you should not take his answer for granted

OpenStudy (xapproachesinfinity):

you should have questioned him how did you do it!

OpenStudy (retireed):

try this approach.... Four books W X Y and Z 34 = W + X +Y 36 = W + X + Z 38 = W + Y + Z 12 = X + Y + Z Do these seem valid to you? by elimination can these be solved ??

OpenStudy (anonymous):

They do seem valid

OpenStudy (retireed):

If I remembered how, solving simultaneous equation on a calculator by using matrices it would be easier. But I am watching Stanford crush Iowa for some reason.

OpenStudy (anonymous):

This question is actually in the practice test of the Hunter College High school entrance exam. Nothing should be complicated. It is a bit hard

OpenStudy (retireed):

Okay at halftime I figured out the calculator and I got W = 28, X = 2, Y = 4 and Z = 6 would be the students reading those books, so the total number of students would be the sum. I checked my answers into the equation and it all checked good.

OpenStudy (anonymous):

Thank you for your help retirEEd

OpenStudy (retireed):

I just can't figure out why Decarr432 answer was correct. Where he took the sum of the books read and divided by 3. ??

OpenStudy (decarr432):

Its because what I did was took four different books I just drew em on paint and put one dash in three different for each student to show them being one person and did that for all of them and it came to 40 unconventional but it worked